# HUMAP: Hierarchical Uniform Manifold Approximation and Projection

### v4

@@ 1 Related Work
Finally, users work with embeddings of increasing size during hierarchical exploration. Thus, it is essential to maintain the mental map between successive projections in order to prevent users from being tricked by geometrical transformations during exploratory analysis. This task, often refered as projection alignment, can be addresses in different ways. A time-varying dataset is projected using dynamic t-SNE [38] to reduce temporal variability that is unimportant to the final layout. This is accomplished by including a term in the cost equation for the t-SNE that regulates the trade-off between alignment and conventional t-SNE optimization. A trade-off parameter is also used in VFF [13] to regulate the alignment of projections in feature fusion tasks. VFF, however, is ineffective for preserving local structures [6]. Finally, Cantareira and Paulovich [6] proposed a general model for projection alignment that incorporates the original cost function of the DR technique and a penalty term applied for alignment. Another important problem that these techniques address is when dealing with streaming applications, where the use of DR techniques for conveying information has grown [MOHEDANOMUNOZ2023120252]. The preservation of mental maps is critical in such scenarios, as users would have to expend a significant amount of cognitive effort tracking the position of data points in the visual space.. Our approach to maintaining the mental map entails directing the movement of data samples that have already been projected at high hierarchy levels and affecting the placement of new data samples. Fig. HUMAP: Hierarchical Uniform Manifold Approximation and Projection illustrates this. The highest hierarchy level (top-level) of HUMAP resembles the lowest hierarchy level and maintains the structures of chosen subsets (B). In contrast to single-level techniques (A), it summarizes the data while providing information on the connections between and within clusters.
@@ 2 Hierarchical Uniform Manifold Approximation and Projection (HUMAP)
The first step in building a hierarchy from bottom to top is to use a kernel function to determine the connection strengths (local affinities) of a $k$-nearest neighbor graph of data points in the high-dimensional space $\mathbb{R}^{m}$ (step A). Then, similar to previous research [36], we employ the Finite Markov Chain (FMC) to identify the most visited nodes, which consist of the landmarks for the higher level (step C). Steps (D) and (E) of the FMC process are used to encode local and global neighborhood information for each landmark as well as to build a neighborhood structure for high hierarchy levels ($>1$). In order to define a new hierarchy level, a new neighborhood graph is created in step (F) using the computed similarity (step G).
@@ 2.2 Dissimilarity definition
The local neighborhood component adapts our approach to preserve more of the local structures (see the pink relationship in Fig. 1). To this end, we use the $k$-nearest neighbors (NH) of each landmark $l^{i}_{u}$ to augment their representation neighborhoods with the $\beta\times|NH(l^{i}_{u})|$ first nearest neighbors of $l^{i}_{u}$, where $\beta$ is a threshold between 00 and $1$. Thus, the greater $\beta$, the more important the $k$-nearest neighbors, and the similarity measure will encode more local relations. HUMAP supports parameter tuning to perform better on neighborhood preservation. The similarity between two landmarks consists of the intersection between their global and local neighborhoods. The local neighborhood for a landmark is in the set of its nearest neighbors, so one can increase it by adjusting the parameter $\beta\in[0,1]$ that adds $\beta|NH(i)|$ neighbors to the local neighborhood—by default, $\beta$ is set to 00. We evaluated HUMAP with different values of $\beta$ and found a clear relationship between local neighborhoods and neighborhood preservation.
Finding these two components results in the representation neighborhood (RNH) for each landmark $l^{i}_{u}$ on level $H^{i}$. Such a neighborhood consists of a sparse matrix $RNH$ of dimensions $|H^{i}|\times|H^{i-1}|$ where $RNH_{u,v}$ is $1$ if $x_{v}^{i-1}\in$ RNH($l_{u}^{i}$) or 00, otherwise. The similarity among landmarks is computed based on the intersection of representation neighborhoods, that is, the shaded area of Fig. 1 (E-F). We compute the intersection with the $RNH$ matrices as follows:
@@
where $M$ is a normalization factor that ensures similarity between 00 and $1$—notice that since $RNH$ is sparse, computing the above matrix multiplication is fast in practice. Note that the intersection between neighborhoods is also used in HSNE to compute the similarity between two landmarks. However, unlike in HSNE, we do not directly employ such a similarity measure to compute the embeddings. Instead, HUMAP extends the neighborhood to encode global information and uses UMAP’s kernel to maintain its properties throughout the hierarchy.
@@ 2.3.2 Mental map preservation
The subsets of data chosen for fine-grained exploration on level $l-1$ during hierarchical exploration must resemble the organization of level $l$, especially for unlabeled datasets. Rauber et al. [38] and Fujiwara et al. [38] also present strategies to deal with mental map preservation in the context of dimensionality reduction, even though the majority of the work on mental map preservation is limited to dynamic graph drawing [50, 24]. In both situations, the goal is to ease the analysis burden during exploration and facilitate smoother embedding layout changes. Fig. 2 shows successive hierarchical projections using HUMAP, whether or not the mental map is preserved. Note that the embedding generated with no mental-map preservation (i.e., $\theta=1$) might deceive the user in thinking that the overall organization of the layout was preserved.
@@
HUMAP uses the coordinates of higher—and already projected—data points on level $l$ to guide the positioning of level $l$-1. When initializing the low-dimensional representation before Stochastic Gradient Descent (SGD) optimization, the coordinates for the projected data points in $l$ are used as starting points together with the remaining data points initialized with Spectral Embedding—these coordinates only move a fraction during SGD optimization. Thus, we also provide this fraction hyperparameter $\theta$ to be tuned according to one’s needs, although we empirically find that $\theta=0.01$ yields embeddings that best preserve the mental map throughout the hierarchy expansion.
@@ 4 Numerical evaluation
We use the following datasets for evaluation. MNIST [23] is a dataset composed of $70,000$ $28\times 28$ pixel grayscale images of handwritten digits classified into ten classes (from 00 to $9$), where each flattened image results in a $784$ dimensional vector. Fashion MNIST [49] (FMNIST) is a dataset composed of $70,000$ $28\times 28$ pixel grayscale images of fashion items (clothing, footwear, and bags) divided into ten classes; each flattened image results in a $784$ dimensional vector. Mammals is a synthetic dataset designed to have four well-separated classes, and it consists of $20,000$ data points described by $72$ dimensions. Embryoid Body (EB) is a single-cell RNA sequencing dataset for embryoid body data generated over $27$ days [31] divided into five periods. For this dataset, we aim to visualize the development of the $31,000$ cells after preprocessing them and using the first 50 principal components to perform the projections. Note that, FMNIST, MNIST, and Mammals datasets are meant to produce cohesive clusters after embedding, while the ideal result for Embryoid Body is to provide an understanding of its continuous structures—this is an attempt to understand how HUMAP compares to different techniques regarding their most known characteristics, that is, to cluster data (HSNE, derived from SNE) and to emphasize continuity (Multiscale PHATE derived from PHATE). The experiments were performed in a computer with the following configuration: Intel(R) Core(TM) i7-8700 CPU @ 3.20 GHz, 32 GB RAM, Ubuntu 64 bits, NVIDIA GeForce GTX 1660 Ti 22 GB.
@@
Figure 6: Runtime execution in seconds using a log10 scale. The boxen-plots show the runtime for $20$ runs of each technique to fit the hierarchy (Hierarchy fit), to embed the top level (Level 2), and to embed the whole dataset (Level 0).
@@ 4.3 DEMaP
We employ the DEMaP metric [30], which calculates the Spearman correlation between geodesic distances on high-dimensional and euclidean distances on low-dimensions, to assess how well the techniques convey manifolds, clusters, and other high-dimensional space structures. On level 2, HUMAP outperformed HSNE, UMAP, and GPU-based HSNE for the MNIST and FMNIST datasets, as shown in Fig. 8. The pair of distributions (HUMAP, UMAP) for the EB, FMNIST datasets on level 2 and for the EB dataset on level 0 are not statistically different after a t-test, but the others are with a p-value of at least $<$ 0.00001 (see Supplementary File - Table 2 for the details). When the entire dataset is embedded for the mammals dataset, HUMAP displays higher values, providing proof that our method is stable across hierarchical levels. While HSNE and Multiscale PHATE produce better results at the top of the hierarchy, they omit key details as one descends hierarchy—the relationship among clusters is lost on the HSNE side. Finally, HUMAP conveys the continuous structures present in the EB dataset even for the top-level embedding while M. PHATE and HSNE CPU show higher DEMaP values at level 0.
@@ 5 Discussion
Another important aspect besides the technical differences among these techniques is related to where they can be employed. Obvious applications are in scenarios where one aims to understand hierachical differences within structures of data, such as structural differences in research papers [Gonzalez2023] and sub-cell types [46]. Finally, DR techniques have been extensively used to monitor the training of machine learning models, which makes hierarchical approaches suitable to summarize information.
@@ Acknowledgements.
The authors wish to express their gratitude to all the reviewers who took the time to review this paper. Your insightful comments and feedback were invaluable and significantly contributed to enhancing the quality and clarity of the manuscript. This work was supported by Fundação de Amparo à Pesquisa (FAPESP) [grant number #2018/17881-3] and the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior (CAPES) [grant number #88887.487331/2020-00].

### v5

@@ 1 Related Work
Finally, users work with embeddings of increasing size during hierarchical exploration. Thus, it is essential to maintain the mental map between successive projections in order to prevent users from being tricked by geometrical transformations during exploratory analysis. This task, often refered as projection alignment, can be addresses in different ways. A time-varying dataset is projected using dynamic t-SNE [40] to reduce temporal variability that is unimportant to the final layout. This is accomplished by including a term in the cost equation for the t-SNE that regulates the trade-off between alignment and conventional t-SNE optimization. A trade-off parameter is also used in VFF [14] to regulate the alignment of projections in feature fusion tasks. VFF, however, is ineffective for preserving local structures [6]. Finally, Cantareira and Paulovich [6] proposed a general model for projection alignment that incorporates the original cost function of the DR technique and a penalty term applied for alignment. Another important problem that these techniques address is when dealing with streaming applications, where the use of DR techniques for conveying information has grown [31]. The preservation of mental maps is critical in such scenarios, as users would have to expend a significant amount of cognitive effort tracking the position of data points in the visual space.. Our approach to maintaining the mental map entails directing the movement of data samples that have already been projected at high hierarchy levels and affecting the placement of new data samples. Fig. HUMAP: Hierarchical Uniform Manifold Approximation and Projection illustrates this. The highest hierarchy level (top-level) of HUMAP resembles the lowest hierarchy level and maintains the structures of chosen subsets (B). In contrast to single-level techniques (A), it summarizes the data while providing information on the connections between and within clusters.
@@ 2 Hierarchical Uniform Manifold Approximation and Projection (HUMAP)
The first step in building a hierarchy from bottom to top is to use a kernel function to determine the connection strengths (local affinities) of a $k$-nearest neighbor graph of data points in the high-dimensional space $\mathbb{R}^{m}$ (step A). Then, similar to previous research [38], we employ the Finite Markov Chain (FMC) to identify the most visited nodes, which consist of the landmarks for the higher level (step C). Steps (D) and (E) of the FMC process are used to encode local and global neighborhood information for each landmark as well as to build a neighborhood structure for high hierarchy levels (>1>1). In order to define a new hierarchy level, a new neighborhood graph is created in step (F) using the computed similarity (step G).
@@ 2.2 Dissimilarity definition
The local neighborhood component adapts our approach to preserve more of the local structures (see the pink relationship in Fig. 1). To this end, we use the $k$-nearest neighbors (NH) of each landmark $l^{i}_{u}$ to augment their representation neighborhoods with the $\beta\times|NH(l^{i}_{u})|$ first nearest neighbors of $l^{i}_{u}$, where $\beta$ is a threshold between 0 and $1$. Thus, the greater $\beta$, the more important the $k$-nearest neighbors, and the similarity measure will encode more local relations. HUMAP supports parameter tuning to perform better on neighborhood preservation. The similarity between two landmarks consists of the intersection between their global and local neighborhoods. The local neighborhood for a landmark is in the set of its nearest neighbors, so one can increase it by adjusting the parameter $\beta\in[0,1]$ that adds $\beta|NH(i)|$ neighbors to the local neighborhood—by default, $\beta$ is set to 0. We evaluated HUMAP with different values of $\beta$ and found a clear relationship between local neighborhoods and neighborhood preservation.
Finding these two components results in the representation neighborhood (RNH) for each landmark $l^{i}_{u}$ on level $H^{i}$. Such a neighborhood consists of a sparse matrix $RNH$ of dimensions $|H^{i}|\times|H^{i-1}|$ where $RNH_{u,v}$ is $1$ if $x_{v}^{i-1}\in$ RNH($l_{u}^{i}$) or 0, otherwise. The similarity among landmarks is computed based on the intersection of representation neighborhoods, that is, the shaded area of Fig. 1 (E-F). We compute the intersection with the $RNH$ matrices as follows:
@@
where $M$ is a normalization factor that ensures similarity between 0 and 11—notice that since $RNH$ is sparse, computing the above matrix multiplication is fast in practice. Note that the intersection between neighborhoods is also used in HSNE to compute the similarity between two landmarks. However, unlike in HSNE, we do not directly employ such a similarity measure to compute the embeddings. Instead, HUMAP extends the neighborhood to encode global information and uses UMAP’s kernel to maintain its properties throughout the hierarchy.
@@ 2.3.2 Mental map preservation
The subsets of data chosen for fine-grained exploration on level l1l-1 during hierarchical exploration must resemble the organization of level $l$, especially for unlabeled datasets. Rauber et al. [40] and Fujiwara et al. [40] also present strategies to deal with mental map preservation in the context of dimensionality reduction, even though the majority of the work on mental map preservation is limited to dynamic graph drawing [52, 25]. In both situations, the goal is to ease the analysis burden during exploration and facilitate smoother embedding layout changes. Fig. 2 shows successive hierarchical projections using HUMAP, whether or not the mental map is preserved. Note that the embedding generated with no mental-map preservation (i.e., $\theta=1$) might deceive the user in thinking that the overall organization of the layout was preserved.
@@
HUMAP uses the coordinates of higher—and already projected—data points on level $l$ to guide the positioning of level ll-1. When initializing the low-dimensional representation before Stochastic Gradient Descent (SGD) optimization, the coordinates for the projected data points in $l$ are used as starting points together with the remaining data points initialized with Spectral Embedding—these coordinates only move a fraction during SGD optimization. Thus, we also provide this fraction hyperparameter $\theta$ to be tuned according to one’s needs, although we empirically find that $\theta=0.01$ yields embeddings that best preserve the mental map throughout the hierarchy expansion.
@@ 4 Numerical evaluation
We use the following datasets for evaluation. MNIST [24] is a dataset composed of $70,000$ $28\times 28$ pixel grayscale images of handwritten digits classified into ten classes (from 0 to $9$), where each flattened image results in a $784$ dimensional vector. Fashion MNIST [51] (FMNIST) is a dataset composed of $70,000$ $28\times 28$ pixel grayscale images of fashion items (clothing, footwear, and bags) divided into ten classes; each flattened image results in a $784$ dimensional vector. Mammals is a synthetic dataset designed to have four well-separated classes, and it consists of $20,000$ data points described by $72$ dimensions. Embryoid Body (EB) is a single-cell RNA sequencing dataset for embryoid body data generated over $27$ days [33] divided into five periods. For this dataset, we aim to visualize the development of the $31,000$ cells after preprocessing them and using the first 50 principal components to perform the projections. Note that, FMNIST, MNIST, and Mammals datasets are meant to produce cohesive clusters after embedding, while the ideal result for Embryoid Body is to provide an understanding of its continuous structures—this is an attempt to understand how HUMAP compares to different techniques regarding their most known characteristics, that is, to cluster data (HSNE, derived from SNE) and to emphasize continuity (Multiscale PHATE derived from PHATE). The experiments were performed in a computer with the following configuration: Intel(R) Core(TM) i7-8700 CPU @ 3.20 GHz, 32 GB RAM, Ubuntu 64 bits, NVIDIA GeForce GTX 1660 Ti 22 GB.
@@
Figure 6: Runtime execution in seconds using a log10 scale. The boxen-plots show the runtime for 2020 runs of each technique to fit the hierarchy (Hierarchy fit), to embed the top level (Level 2), and to embed the whole dataset (Level 0).
@@ 4.3 DEMaP
We employ the DEMaP metric [32], which calculates the Spearman correlation between geodesic distances on high-dimensional and euclidean distances on low-dimensions, to assess how well the techniques convey manifolds, clusters, and other high-dimensional space structures. On level 2, HUMAP outperformed HSNE, UMAP, and GPU-based HSNE for the MNIST and FMNIST datasets, as shown in Fig. 8. The pair of distributions (HUMAP, UMAP) for the EB, FMNIST datasets on level 2 and for the EB dataset on level 0 are not statistically different after a t-test, but the others are with a p-value of at least << 0.00001 (see Supplementary File - Table 2 for the details). When the entire dataset is embedded for the mammals dataset, HUMAP displays higher values, providing proof that our method is stable across hierarchical levels. While HSNE and Multiscale PHATE produce better results at the top of the hierarchy, they omit key details as one descends hierarchy—the relationship among clusters is lost on the HSNE side. Finally, HUMAP conveys the continuous structures present in the EB dataset even for the top-level embedding while M. PHATE and HSNE CPU show higher DEMaP values at level 0.
@@ 5 Discussion
Another important aspect besides the technical differences among these techniques is related to where they can be employed. Obvious applications are in scenarios where one aims to understand hierachical differences within structures of data, such as structural differences in research papers [13] and sub-cell types [48]. Finally, DR techniques have been extensively used to monitor the training of machine learning models, which makes hierarchical approaches suitable to summarize information.
@@ Acknowledgements.
The authors wish to express their gratitude to all the reviewers who took the time to review this paper. Your insightful comments and feedback were invaluable and significantly contributed to enhancing the quality and clarity of the manuscript. This work was supported by Fundação de Amparo à Pesquisa (FAPESP) [grant number #2018/17881-3] and the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior (CAPES) [grant number #88887.487331/2020-$0$].